The average number of miles driven on a full tank of gas in a certain model car before its low-fuel light comes on is 331. AsThe answers are in red. Please explain all, esp part d!
Answer #1

a) Let X be the number of miles driven by the car before the low fuel light turns on

XN(331, 342

then

-331 N(0, 1) 34

349-331 Р(X < 349) — P( < 34

= P( z < 0.53)

= 0.7019 (from z table)

(b)

273 331 P(X 273) P(z> 34

= P( z > -1.71)

= 0.9564 (from z table)

(c)

271 331 291 331 P(271 < X< 291) P( くく 34 34

= P( -1.76 < z < -1.18 )

= P( -1.76 < z < 0) - P( -1.18 < z < 0)

= 0.4608 - 0.3810 (from z table)

= 0.0798

(d) Using continuity correction

P( X=303) = P( 302.5<X < 303.5)

= P( -0.84 < z < -0.81)

= 0.2995 - 0.2910   (from z table)

= 0.0085

X-N(331, 34)

X - 331 E N (0,1)

349-331 P(X <349) = P(Z <= 34

273 - 331 P(X > 273) = P(Z >= 34

271 - 331 P(271 < X < 291) = P(- 291 - 331 34 34